if triangle ABC is an isosceles triangle, aangle B is the vertex angle, AB=6x+3 BC=8x-1 and AC=10x-10 find x and the measures of each side of the triangle
Answer by venugopalramana(3286)

if triangle ABC is an isosceles triangle, angle B is the vertex angle, AB=6x+3 BC=8x-1 and AC=10x-10 find x and the measures of each side of the triangle
ABC is an isosceles triangle, angle B is the vertex angle..hence BA=BC..or..AB=BC...hence...
6x+3=8x-1
8x-6x=3+1=4
2x=4
x=4/2=2
so the 3 sides are
AB=6x+3=6*2+3=12+3=15
BC=8x-1=8*2-1=16-1=15
AC=10x-10=10*2-10=20-10=10